A car is speeding up and has an instantaneous velocity of 1.0 m/s in the +x-direction when a stopwatch reads 10.0 s. It has a constant acceleration of 2.0 m/s2 in the x-direction. (a) What change in speed occurs between t = 10.0 s and t = 12.0 s? (b) What is the speed when the stopwatch reads 12.0 s?


1
Expert's answer
2022-03-13T18:49:19-0400

a) The change in speed is given as follows, according to the kinematic equation:



"\Delta v = a\Delta t"

where "a= 2m\/s^2" is the acceleration, and "\Delta t = 12s-10s = 2s" is the time interval. Thus, obtain:



"\Delta v = 2m/s^2\cdot 2s = 4m/s"

b) The speed at 12s is:



"v = v_0 + \Delta v"

where "v_0 = 1m\/s" is the initial speed. Thus, have:



"v = 1m\/s + 4m\/s = 5m\/s"

Answer. a) 4m/s, b) 5m/s.